# A Neural Network Enhanced Born Approximation for Inverse Scattering

### v2

@@ 5.2 Noise robustness
In contrast, the CNNB model exhibits remarkable robustness to noise, consistently producing accurate reconstructions with minimal artifacts even under high noise conditions. The regularized inversion incorporated into the CNNB algorithm effectively controls noise amplification and enables the optimized selection of regularization parameters. This capability demonstrates the model’s ability to reconstruct scatterers accurately in noisy environments. Figure 6 illustrates the comparative performance of the three models at different noise levels.
[…]
5.3 Absorption
Our CNNs are trained on data from dissipation free media ($\Im(\eta)=0$). In many practical applications, the media exhibits varying degrees of dissipation or loss [12], characterized by $\Im(\eta)>0$. Since we did not train using far field data from dissipative scatterers, it is important to determine if we can predict the real part of $\eta$ even when the scattering media are slightly dissipative. The objective of this experiment is to evaluate the CNN models’ ability to accurately reconstruct $\Re(\eta(\textbf{x}))$ despite the presence of absorption in the true scatterer.
To investigate such scenarios, we consider the case in which $\eta$ inside the circle is perturbed by an unknown, small absorption term. Specifically, we use the following complex-valued $\eta$ to generate far field test data:
$\eta^{\alpha}\mathrel{\mathop{:}}=\eta_{0}+\mathrm{i}\mkern 1.0mu\alpha$ (23)
[…]
The results are shown in Figure 7. We observe that the CNNB architecture can effectively handle the exponential decay in the total field caused by absorption in the scatterer, making it a promising candidate for practical applications where absorption effects, though small, are present. In contrast, the BCNN and CNN networks are less stable to absorption and fail to preserve the shape of the scatterer for $\alpha=10$.
[…]
@@ 5.4 Increased scatterer complexity
We next evaluate the generalization capabilities of the CNN models when applied to more complex out-of-distribution scatterers. The scatterers are as follows:
  1. 1.
    A U-shaped scatterer chosen to resonate with the incident field (see top left panel in Figure 8). As we shall see, for this shape, the opening of the U is difficult to image using the inverse Born method.
  2. 2.
    A high-contrast annulus or ring inside a circle with an elliptic inclusion as shown in Figure 9. The goal with this example is to test if the CNN based schemes can detect the inner wall, even though they were trained on scattering by simple disks.
  3. 3.
    In potential biomedical applications, it is desirable to determine the refractive index of structures within other structures. In this example, we construct a scatterer akin to the Shepp-Logan phantom. High-contrast scatterers are placed within a high-contrast ring as shown in the top left hand panel of Figure  10. Restrictions on our mesh generator for the forward problem prevented us using the Shepp-Logan phantom itself. We also note that the values of $\eta$ used here are not the same as for the real Shepp-Logan phantom.
  4. 4.
    The CNN based methods are trained on smooth scatterers. Our last test uses a scatterer made of rectangles and an inner low contrast region (see top left panel in Figure 11).
The results of running the various reconstruction algorithms are summarized in Table 3 and in Figures 811 and are discussed next:
@@
Table 3: Relative $\ell^{2}$ error in the reconstruction of $\mu$ as percentages for the inverse Born approximation and the three CNN models for the complex scatterers in Section 5.4. The best reconstruction for each problem is in boldface.
Relative $\ell^{2}$ Error (%)
Method U Ring Shepp Rectangles
Born ($\gamma=0.1$) 92.72 106.26 104.44 87.68
Born ($\gamma=1$) 88.95 87.49 92.76 81.60
CNN 67.63 44.41 63.42 43.66
CNNB 60.53 41.23 60.03 39.28
BCNN 66.28 39.76 60.57 43.03
  1. 1.
    The results of reconstructing the high-contrast U-shaped resonant structure are shown in Figure  8 and Table 3. The inverse Born scheme exhibits artifacts and does not clearly show the U. While all three CNN models greatly improve on the Born approximation, the CNNB model preserves the shape of the U with minimal background artifacts. Meanwhile, the BCNN model constructs the scatterer with distortion in the surrounding media and the CNN model is blurred. The relative error in $\ell^{1}$ or $\ell^{2}$ conform that CNNB is best in this case.
  2. 2.
    Our second example, the high contrast ring scatterer, tests if it is possible to observe the inner structure of an object. As shown in Figure 9, the direct use of the inverse Born scheme does not reveal the inner ellipse. All three CNN models improve of the inverse Born, and perform relatively similarly in obtaining the shape. In this case, the BCNN model best approximates the true value of $\eta$ in the annulus as shown in Table 3.
  3. 3.
    Concerning the modified Shepp-Logan phantom, we see from Figure 10 and Table 3 that with appropriate regularization, the inverse Born scheme correctly images (though under-approximates) $\eta$ in the outer boundary of the scatterer, but fails to image structures inside. All CNN based methods improve over the inverse Born approximation, though the BCNN model well-approximates both the outer ring and the separation of the three internal scatterers, CNNB gives the best quantitative reconstruction.
  4. 4.
    We finally consider the shape constructed from several rectangles, possessing both sharp corners and an interior of air as shown in Figure 11 and Table 3. We see that the inner inclusion is invisible to the Born approximation, and the high contrast block is not well approximated. All three CNN models perform comparatively similarly, but the CNNB and BCNN models better approximate the shape of the internal region in comparison to the CNN and Born models. Quantitatively CNNB performs best. The presence of corners does not badly impact the CNN based models.
Overall we have demonstrated good generalization to a variety of scatterers not obviously connected to the training data. These results hint that training on simple shapes like circles is a successful strategy.
@@
Figure 8: The reconstruction of a high-contrast, U-shaped resonance structure by the inverse Born approximation and the there CNN based schemes. All three CNN based schemes result in much clearer reconstructions with CNNB having less background noise than BCNN or CNN.
@@
Figure 9: The reconstruction of a high-contrast annular scatterer. All three CNN based schemes are markedly better than inverse Born alone. With the CNN approach, the interior of the annulus can be reconstructed.
@@
Figure 10: The reconstruction of a simplified Shepp-Logan-like phantom. Here the scattering ellipses are surrounded by a thin high contrast annulus of varying thickness. The Born approximation can detect this ring, but fails to reconstruct inside. All three CNN based schemes detect the three internal scatterers.
@@
Figure 11: The reconstruction of a structure with sharp corners. The inner void is not seen in the Born reconstructions but is detected by the CNN schemes with CNNB and BCNN producing better reconstructions than CNN alone.
@@ 6 Conclusion
Our examples demonstrate that combining CNNs with the Born approximation can extend the applicability of the method beyond the weak scattering limit. All three CNN models can reveal the inner structure of high contrast objects better than the Born approximation. Moreover, a CNN combined with the inverse Born model always improves scatterer reconstruction in comparison to a pure CNN model. Because of it’s stability to noise and optimality for almost every test, we prefer the CNNB approach.
It is notable that we elected to train using simple shapes (disks), but the trained CNN models successfully generalize to more exotic cases.
Much more work needs to be done to extend this demonstration to a realistic biomedical problem like ultrasound tomography. For example, the current implementation of the regularized inverse Born approximate uses simple dense linear algebra to evaluate the necessary inverse. Fast methods should be used, such as the specialized neural network of [26] or the low-rank method of [25]. In addition, for specific applications, other simple training data could be considered (for example, if the goal is to image a network of blood vessels).
A particularly interesting direction for further work is to apply the method to more exotic measurement scenarios. For example, the case where both transmitters and receivers are on one side of the object, or when there is missing data from certain angular sectors. Moreover, additional work could include extending the applicability of the model to cases in which the wave number $k$ varies (i.e. multi-frequency data) or when the scattering medium is defined by a function that is not piecewise-constant.
@@ Acknowledgments
The research of P.M. is partially supported by the US AFOSR under grant number FA9550-23-1-0256. The research of T.L. is partially supported by the Research Council of Finland via the Finnish Center of Excellence of Inverse Modeling and Imaging, the research project 321761, and the Flagship of Advanced Mathematics for Sensing Imaging and Modeling grant 358944. The research of A.D. is supported by the University of Delaware Undergraduate Research Program. This research was supported in part through the use of Information Technologies (IT) resources at the University of Delaware, specifically the high-performance computing resources. The authors also acknowledge the CSC – IT Center for Science, Finland, for generously sharing their computational resources.

### v3

@@ 5.3 Absorption
The results are shown in Table 4. We observe that, judged using the $\ell^{2}$ norm, the NIO architecture performs best under weak absorption for $\alpha=0.1$ in the scatterer. CNNB also tolerates small absorption and is the best of the CNN based models. However, all models deteriorate for higher magnitudes of absorption ($\alpha=1,10$) and BCR-Net performs optimally in the $\ell^{2}$ in this regime.
Fig. 9 shows the results for one reconstruction at various absorption levels of the ground truth scatterer in Fig. 8. In this case, NIO does not successfully reconstruct the smallest scatterer. BCR-Net places this scatterer further from the large scatterer and hallucinates by predicting a fourth scatterer for $\alpha=1,10$.
Figure 9: Here we show the refractive index estimated by each model with data from an absorbing scatterer corresponding to the ground truth scatterer ($\eta_{0}=1.8$) in Fig. 8. For larger absorption values, the CNN based models tend to under-approximate with $\eta_{0}\approx 1.3$ and produce blurry artifacts surrounding the scatterers. Meanwhile, NIO over-estimates with $\eta_{0}\approx 2.1$ and fails to distinguish the smaller scatterer, while BCR-Net predicts a spurious fourth scatterer.
@@ 5.4 Increased scatterer complexity
We next evaluate the generalization capabilities of the models when applied to more complex out-of-distribution scatterers. For each scatterer type, we generate 100 instances with random type-specific variations. The performance of the various reconstruction algorithms are summarized in Table 5 and sample instances are displayed in Fig. 1013.
@@
Table 5: Average relative $\ell^{2}$ and $\ell^{1}$ errors (%) in the reconstruction of the contrast $\mu$ on 100 randomized samples of the complex scatterers discussed in Section 2. The best reconstruction for each problem is in boldface.
U Ring
Method $\ell^{2}$ (%) $\ell^{1}$ (%) Method $\ell^{2}$ (%) $\ell^{1}$ (%)
Born ($\gamma=1$) 69.6645 119.4050 Born ($\gamma=1$) 109.3183 116.4062
CNN 55.8611 110.5492 CNN 64.7712 65.2693
NIO 77.8722 109.2033 NIO 94.3889 98.3254
CNNB 53.5359 98.3680 CNNB 60.5713 63.8227
BCNN 53.9426 130.5083 BCNN 74.5521 82.6612
BCR-Net 80.3159 105.6085 BCR-Net 73.9569 74.0454
Shepp-Logan Rectangles
Method $\ell^{2}$ (%) $\ell^{1}$ (%) Method $\ell^{2}$ (%) $\ell^{1}$ (%)
Born ($\gamma=1$) 97.9434 138.3105 Born ($\gamma=1$) 89.5299 118.3325
CNN 59.2884 84.0929 CNN 42.6187 58.0487
NIO 75.05212 91.7023 NIO 57.4389 59.3596
CNNB 58.5902 87.1067 CNNB 42.0379 55.1703
BCNN 59.0327 102.3163 BCNN 45.9519 75.3718
BCR-Net 78.3657 87.3097 BCR-Net 73.8079 75.08584
5.4.1 Resonance Structure
Multiple scattering induced by resonance structures can often produce challenging far field patterns that prevent accurate inversion or reconstruction with the Born approximation. We test this using a U-shaped scatterer chosen to resonate with the incident field, as shown in Fig.10. Here, we fix $\eta=1.8$ (strong scattering).
To generate random test data, the location, size, and rotation of each sample structure are chosen randomly from the uniform distributions $\mathcal{U}(-0.9,0.9)$, $\mathcal{U}(0.05,0.2)$, and $\mathcal{U}(0,2\pi)$, respectively (but keeping the U within $[-1,1]^{2}$). In this case, results in Table 5 show that CNNB outperforms the other methods. To understand this, Fig. 10 shows a particular realization of the inversion. In this case, the inverse Born method does not approximate the the magnitude of $\eta$ well. While all three CNN models improve on the Born approximation, the CNNB model preserves the shape of the U with minimal background speckle. Meanwhile, BCNN constructs the scatterer with distortion in the surrounding media and the CNN model is blurred. NIO produces blurred artifacts that do not resemble the U-shaped structure, while BCR-Net fails to find the scatterer.
@@ 5.4.1 Resonance Structure
Figure 10: The reconstruction of a high-contrast, U-shaped resonance structure by the inverse Born approximation and the NN based schemes. All three CNN based schemes result in much clearer reconstructions with CNNB having less background noise than BCNN or CNN. The NIO approach fails at constructing both the interior and the walls of the structure.
5.4.2 Ring
The goal with this example is to test if the NN based schemes can detect an inner wall, even though they were trained on scattering by simple disks. We construct an annulus or ring inside a circle with an elliptic inclusion, as shown in Fig.  11. For each sample, the value of $\eta$ in each component are chosen randomly from the uniform distribution $\mathcal{U}(1.1,1.8)$. Averaged results are shown in Table 5 showing CNNB outperforms the other methods.
For a particular choice of the model parameters, Fig. 11 shows that the direct use of the inverse Born scheme may not reveal the inner ellipse. In this case all three CNN models improve the inverse Born approximation and perform relatively similarly in obtaining the shape. Meanwhile, NIO and BCR-Net do not produce an accurate representation of the scatterer.
@@ 5.4.2 Ring
Figure 11: The reconstruction of a high-contrast annular scatterer. All three CNN based schemes are markedly better than inverse Born alone at representing the interior ellipses. However, the outer annular structure is poorly estimated. The NIO and BCR-Net approaches fail to detect the outer annulus.
5.4.3 Shepp-Logan Phantom
In potential biomedical applications, it is desirable to determine the refractive index of structures within other structures. In this example, we construct a scatterer akin to the Shepp-Logan phantom [38], as shown in Fig. 12. Restrictions on our mesh generator for the forward problem prevented us from using the Shepp-Logan phantom itself. To generate statistical data, for each sample the value of $\eta$ in each component is chosen randomly from the uniform distribution $\mathcal{U}(1.1,1.8)$ so that the values of $\eta$ used here are not the same as for the real Shepp-Logan phantom. Table 5 summarizes the statistical results showing that CNNB is best in the $\ell^{2}$ norm but CNN outperforms it in the $\ell^{1}$ norm.
In the particular case shown in Fig. 12, we see that, with appropriate regularization, the inverse Born scheme correctly images (though under-approximates) $\eta$ in the outer boundary of the scatterer, but fails to image structures inside. All CNN based methods improve over the inverse Born approximation, though the BCNN model well-approximates both the outer ring and the separation of the three internal scatterers, CNNB gives a good quantitative reconstruction. Again, we observe that NIO and BCR-Net fail to produce an accurate representation of the phantom.
@@ 5.4.3 Shepp-Logan Phantom
Figure 12: The reconstruction of a simplified Shepp-Logan-like phantom. Here the scattering ellipses are surrounded by a thin high contrast annulus of varying thickness. The Born approximation can detect this ring, but fails to reconstruct inside. All three CNN based schemes detect the three internal scatterers, though CNNB produces the sharpest image. In contrast, NIO and BCR-Net do not produce any meaningful representation of the ground truth. BCR-Net also does not capture the overall structure at all.
5.4.4 Rectangles
The NN based methods are trained on smooth scatterers. Our last test uses a scatterer made of rectangles and an inner void of air, as shown in Fig. 13. For each testing sample, the value of $\eta$ in the left and right pieces are chosen randomly from the uniform distribution $\mathcal{U}(1.1,1.8)$. Table 5 shows that CNNB is, on average, the best scheme.
For the particular realization shown in in Fig. 13 we see that the inner inclusion is invisible to the Born approximation, and the high contrast block is not well approximated. All three CNN models perform comparatively similarly, but the CNNB and BCNN models better approximate the shape of the internal region. While NIO is able to detect the interior void, it performs poorly in approximating the walls of the scatterer, while BCR-Net does not detect the void. In general, the presence of corners in this case does not badly impact the CNN based models.
@@ 5.4.4 Rectangles
Figure 13: The reconstruction of a structure with sharp corners. The inner void is not seen in the Born reconstructions but is detected by the CNN schemes with CNNB and BCNN producing better reconstructions than CNN alone. NIO produces a rough inner void but fails to handle the sharp corners of the scatterer walls.
5.4.5 Discussion
Overall, we have demonstrated good generalization to a variety of scatterers not obviously connected to the training data. These results hint that training on simple shapes like circles is a successful strategy. Moreover, we observe the quantitative success of CNNB in extending to out-of-distribution scatterers through its optimality in $\ell^{2}$ and in $\ell^{1}$ (except for Shepp-Logan), as displayed in Table 5.
While we do not have a definitive explanation for why the CNN based schemes perform well in out-of-distribution scenarios, we conjecture that the local properties of CNNs, through their convolution blocks, may promote added structure [16]. In addition, the inclusion of the Born approximation in CNNB and BCNN introduces useful physical information that biases the learning process toward physically consistent reconstructions. This may improve generalization by implicitly regularizing the solution space, even beyond the weak-scattering regime.
@@ 6 Conclusion
Our tests demonstrate that combining CNNs with the Born approximation can extend the applicability of the Born approximation beyond the weak scattering limit. Moreover, a CNN combined with the inverse Born model always improves scatterer reconstruction in comparison to a pure CNN model.
Comparing to state of the art networks, CNNB performs similarly to NIO and BCR-Net on in-distribution statistical tests, although we have noted that in some cases NIO and BCR-Net fail to detect low contrast and small scatterers in close proximity to larger and higher-contrast scatterers. On noisy data, BCR-Net outperforms all the other networks although CNNB is also stable to low noise but BCNN is less stable. Both NIO and BCR-Net sometimes hallucinate at high noise levels.
When inverting data from absorbing scatterers, NIO outperforms BCR-Net for low absorption. It is noticeable that, in some realizations of the test data, BCR-Net incorrectly predicts an additional scatterer.
The CNN and Born based inversion methods excel at reconstructing the more complex out of distribution scatterers. In this case CNNB is always best in the $\ell^{2}$ norm, and because of this and its stability to noise, we prefer the CNNB approach compared to CNN or BCNN. These findings highlight that incorporating physics-based structure such as the Born approximation into learning models can improve both reconstruction accuracy and the model’s ability to generalize to different data regimes.
Much more work needs to be done to extend this demonstration to a realistic biomedical problem like ultrasound tomography. Fast methods should be used to compute the inverse Born approximation, such as the specialized NN of [46] or the low-rank method of [45]. Further, other inversion algorithms, such as an $L^{1}$ (LASSO) or total-variation based regularization of the Born step, could produce sharper results from CNNB at the cost of longer computing times. In addition, for specific applications, other simple training data could be considered (for example, if the goal is to image a network of blood vessels).
Another particularly interesting direction for further work is to apply the method to more exotic measurement scenarios. For example, the case where both transmitters and receivers are on one side of the object, or when there is missing data from certain angular sectors. Moreover, additional work could include extending the applicability of the model to cases in which the wavenumber $k$ varies (i.e. multi-frequency data) or when the scattering medium is defined by a function that is not piecewise-constant.
@@ Acknowledgments
This research was supported in part through the use of the DARWIN computing system: DARWIN – A Resource for Computational and Data-intensive Research at the University of Delaware and in the Delaware Region, which is supported by NSF under Grant Number: 1919839, Rudolf Eigenmann, Benjamin E. Bagozzi, Arthi Jayaraman, William Totten, and Cathy H. Wu, University of Delaware, 2021, URL:https://udspace.udel.edu/handle/19716/29071. The authors acknowledge the CSC – IT Center for Science, Finland, for generously sharing their computational resources. We thank Dr. Yuwei Fan and Professor Lexing Ying for making the source code for BCR-Net available to us. We also thank Professor Ke Chen of the University of Delaware for sharing his computational resources. Finally, we thank the anonymous referees for their very valuable suggestions.